OpenStudy (anonymous):

A large tank contains 100 liters of a salt solution which has a concentration of 0.25 kilograms per liter. Pure water is added to the solution at a rate of 80 liters per minute. At the same time, the well-mixed solution drains from the tank at 80 liters per minute. Find the amount S of the salt in the solution as a function of t: kilograms At what time is there only 12 kilograms of salt remaining in the tank? t= minutes

OpenStudy (anonymous):

Hi dichalao, Do you still need a solution?

OpenStudy (anonymous):

Yes! Or just give me some clues first :)

OpenStudy (anonymous):

What i have tried is to set up the first differential equation as : dS/dt= 25/(100+80t)(1-80t)

OpenStudy (anonymous):

You need to consider the physics of what's going on in a little bit of time,\[\delta t\]

OpenStudy (anonymous):

and you need to consider what's actually changing. You're told 80 L/min enters the tank, and 80 L/min leaves the tank, so volume's the same. What's changing is the concentration, or even better, the mass, of salt in the solution.

OpenStudy (anonymous):

You need to come up with an expression for the change in mass of salt in the tank as a function of time, then use that to get an expression for concentration and solve the system.

OpenStudy (anonymous):

Yea i see the solution is diluting

OpenStudy (anonymous):

So, let's consider what happens in that little bit of time, delta_t. No salt is entering the system, since what's coming in is pure water. Salt *is* leaving the system in the outflow. I said we need to consider mass, so, let the mass at time t be\[m(t)\] and that at time t+delta_t be\[m(t+\delta t)\]The difference between the two will be\[\delta m = m( t + \delta t) - m(t)\]

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

Hi Lok!!

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

The change in mass is the negative of the amount leaving (because these delta quantities are always read (final)-(initial), and since you have less at the end, the change will be negative). That is,\[\delta m = -Q \delta t C(t)\]where Q is the flow rate (liters/min) and C(t) is concentration at time t (this is why delta_t needs to be small, otherwise C(t) won't be valid).

OpenStudy (anonymous):

Notice the units on the left-hand side and the right-hand side. LHS is (mass), and right-hand side is (volume)/(time) x (time) x (mass)/(volume). The units equate.

OpenStudy (anonymous):

So what we have now is\[\delta m = m(t+ \delta t)-m(t)\]\[-Q \delta t C(t)=m(t+\delta t)-m(t)\]Divide both sides by the constant volume in the container, and by delta_t to get\[-\frac{QC(t)}{V}=\frac{C(t+\delta t)-C(t)}{\delta t}\]

OpenStudy (anonymous):

because \[C(t) = \frac{m(t)}{V}, C(t+\delta t)=\frac{m(t+\delta t)}{V}\]

OpenStudy (anonymous):

That expression above (above what I *just* wrote) should look familiar. Does it?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Sorry, my computer's acting up.

OpenStudy (anonymous):

You can now take the limit as delta_t approached zero on both sides to get,

OpenStudy (anonymous):

\[\lim_{\delta t \rightarrow 0}-\frac{QC(t)}{V}=\lim_{\delta t \rightarrow 0}\frac{C(t+ \delta t)-C(t)}{\delta t}\]\[\rightarrow -\frac{QC(t)}{V}=\frac{dC}{dt}\](the left-hand side is independent of delta_t, so it's just a constant in terms of taking the limit).

OpenStudy (anonymous):

You now have a separable differential equation:\[\frac{dC}{C}=-\frac{Q}{V}dt \rightarrow \int\limits_{}{}\frac{dC}{C}=\int\limits_{}{}-\frac{Q}{V}dt\]

OpenStudy (anonymous):

Thanks i can take it from here on

OpenStudy (anonymous):

\[\log C=-\frac{Q}{V}t+c\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I have one more question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Suppose taht a_n>0, and lim n->infinity n^2a_n=0. Prove that sigma a_n converges

OpenStudy (anonymous):

Since lim n^2a_n->0, so i assume n^2a_n behave like 1/n^p where p >0, if so a_n should be have like (1/n^(2+p), a_n is a p-series, with p >1, so a_n converges. Do u think this approach is reasonable?

OpenStudy (anonymous):

You might have to leave it with me. I just got up and need to get ready for uni/work. I'll get back to you.

OpenStudy (anonymous):

Sure :) Have a good one :)

OpenStudy (anonymous):

Before I go, I want to confirm the sequence is\[n^2a_n\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have a question regarding C(t) is concentration at time t (this is why delta_t needs to be small, otherwise C(t) won't be valid).

OpenStudy (anonymous):

Do you mean C(t)*delta_t wont be valid? Or just C(t)

OpenStudy (anonymous):

Well, C(t) is the concentration at time t, but that's always changing. We're looking at times t and t+delta_t, but only using C(t) to approximate the amount of mass that's leaving. This is valid in the limit, but if you leave everything as a finite difference (which is what you have before taking the limit), what you're really saying is\[-QC(t)/V \approx m(t+ \delta t)-m(t)\]

OpenStudy (anonymous):

But because our increment in time was small, the CHANGE in C(t) over that time is assumed very small too.

OpenStudy (anonymous):

So then all you need to do to find the mass is multiply it by volume flow rate.

OpenStudy (anonymous):

I don't know how well this is being explained. We could have used differentials instead of delta notation, but I think understanding the approximations before taking limits is useful.

OpenStudy (anonymous):

This is a good explanation, i understand everything you did

OpenStudy (anonymous):

-Q is -80L/min?

OpenStudy (anonymous):

No, Q=-80L/min

OpenStudy (anonymous):

I took Q as the magnitude of flow, that's why I introduced the minus sign when setting up the equation.

OpenStudy (anonymous):

You should end up with\[C(t)=\frac{1}{4}e^{-\frac{4}{5}(\min^{-1})t}kg/L\]

OpenStudy (anonymous):

which makes sense, since as t approaches infinity, C(t) approaches zero...the salt is being drained away.

OpenStudy (anonymous):

and at time 0, C(0) = 1/4 kg, which is what you've stated as an initial condition.

OpenStudy (anonymous):

1/4 kg/L, I mean.

OpenStudy (anonymous):

if i neeed to find the amount S of the salt in the solution as a function of t: i just need to multiply both side by 100? since V is constant?

OpenStudy (anonymous):

right side*

OpenStudy (anonymous):

Yes, (concentration) x (volume) = (mass) here.

OpenStudy (anonymous):

got it~

OpenStudy (anonymous):

If the volume flow rates were different either side, you'd have a very different situation to look at.

OpenStudy (anonymous):

lol it will be super complicated?

OpenStudy (anonymous):

Kind of...

OpenStudy (anonymous):

where are you from btw

OpenStudy (anonymous):

So, in these types of problems, look at what's actually changing, and try to come up with an expression that focuses on that. Try to formulate something you can turn into a differential equation in the end.

OpenStudy (anonymous):

UK citizen in Sydney.

OpenStudy (anonymous):

Where are you from? Are you male or female?

OpenStudy (anonymous):

I am from US, a kid

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

So is this school stuff?

OpenStudy (anonymous):

yea, i am learning sequence/series and my friend is learning 1st degree differential equation

OpenStudy (anonymous):

You are really good at problem solving

OpenStudy (anonymous):

Is this for the international baccalaureate?

OpenStudy (anonymous):

haha, thanks

OpenStudy (anonymous):

international baccalaureate?

OpenStudy (anonymous):

hey, I just found the answer to your other problem

OpenStudy (anonymous):

haha do you have tips on how to approach a solution?

OpenStudy (anonymous):

AMAZING !!

OpenStudy (anonymous):

the stuff you're studying is covered in the international baccalaureate - it's an international diploma that can be used to get into universities around the world.

OpenStudy (anonymous):

actually, hang on, I'll need to check it. I was looking to apply the Dirichlet test with a series 1/n^2...but I want to see if I can get around one of the assumptions

OpenStudy (anonymous):

No it is just normal school work

OpenStudy (anonymous):

I'll think about tips on how to approach a solution and get back to you. Really, all you have to do is determine what you're being asked to find, put down the information you have before you and build a bridge.

OpenStudy (anonymous):

are you told that \[a_n \ge a_{n+1}>0?\]

OpenStudy (anonymous):

No, what i posted is everything given

OpenStudy (anonymous):

i wouldn't have thought about setting -Qdelta_t(Ct)=m(delta_t+t)-m(t) haha

OpenStudy (anonymous):

Did you ask yourself what was actually changing in the system?

OpenStudy (anonymous):

Concentration's harder to come to terms with than mass, since concentration is a ratio of quantities, whereas mass is something we have an intuitive feel for.

OpenStudy (anonymous):

experience helps too

OpenStudy (anonymous):

yea i know that the amount of salt in the container is lessening and the amount of salts going out is also lessening

OpenStudy (anonymous):

How old are you ~

OpenStudy (anonymous):

if you dont mind :)

OpenStudy (anonymous):

20s

OpenStudy (anonymous):

You are a genius i guess ..

OpenStudy (anonymous):

like I said, experience helps

OpenStudy (anonymous):

You should pass some experiences to me haha

OpenStudy (anonymous):

I have been! =D

OpenStudy (anonymous):

LOL True.. Are you still trying to find the solution to the problem?

OpenStudy (anonymous):

yes, but I should be showering...I have an idea, but not much time to check it out.

OpenStudy (anonymous):

take your time buddy, i have to go to class right now.. Thank you for helping out :) see you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Damn, I figured it out just after you left. You can use the limit comparison test.

OpenStudy (anonymous):

The version we need says, If \[\lim_{n \rightarrow \infty} \frac{a_n}{b_n}=0\]then\[\sum_{n=0}^{\infty}b_n\]convergent implies\[\sum_{n=0}^{\infty}a_n\]is also convergent.

OpenStudy (anonymous):

Now, you need to show that the sum of a_n is convergent, and we can consider for out 'b' series, \[\sum_{n=0}^{\infty}b_n=\sum_{n=0}^{\infty}\frac{1}{n^2}\]which is convergent (p-series, p=2>1). Therefore, taking the ratio of a_n and b_n and forming the limit gives,\[\lim_{n \rightarrow \infty}\frac{a_n}{b_n}=\lim_{n \rightarrow \infty}\frac{a_n}{1/n^2}=\lim_{n \rightarrow \infty}n^2a_n=0\]by construction (i.e. by the setup of the problem). By the theorem above, your series is convergent.

OpenStudy (anonymous):

HAHA got it !!!!

OpenStudy (anonymous):

Yeah. Whenever I see a polynomial in the variable, I think ratio test, limit comparison or Dirichlet first.

OpenStudy (anonymous):

A large tank contains 450 liters of a salt solution which has a concentration of 0.65 kilograms per liter. A salt solution which has a concentration of 0.8 kilograms per liter is added to the solution at a rate of 20 liters per minute. At the same time, the solution drains from the tank at 20 liters per minute.

OpenStudy (anonymous):

It's going to be similar. Volume in the tank won't change again, which is useful. Try to form a mass balance like we did before.